Stephanie was injured in a car accident and was rushed to the emergency room. She received stitches for a facial wound and treatment for a broken finger. Under Stephanie’s PPO plan, emergency room care at a network hospital is 70 percent covered after the member has met a $300 annual deductible. Assume that Stephanie went to a hospital within her PPO network. Her total emergency room bill was $840.

(a) What amount did Stephanie have to pay? (Do not round intermediate calculations.)

Insured's cost=

(b) What amount did the PPO cover? (Do not round intermediate calculations.)

Insurance payment=   
Answer #1

Coinsurance Expenses= Total Expenses - Deductible Expenses

=840-300 =$540

a) Insured's Cost= [Deductible +(1-.70)* Coinsurance expenses]

=[300+0.30*540]

= 300+162 =$462

b) Insurance Payment= Insurance Percent * Coinsurance Expenses

= 0.70*540 =$378

If you have any doubt please write in comment section and do not forget to hit like button.